1. the difference of five times a number and 5 results in twice the difference of three times the number and 3.
the unknown number is?

Answer by ikleyn(53937)

.
According to the condition,  an equation for the unknown number  x  is

5x - 5 = 2*(3x - 3).

Solve it:

5x - 5 = 6x - 6

5x - 6x = -6 + 5,

-x = -1,

x = 1.